### v1

# Rarefied neutral flow in orificed hollow cathodes

Abstract
A parametric study is conducted to quantify the effect of the keeper electrode geometry on the neutral flow quantities (pressure, density, temperature) within orificed hollow cathodes, prior to cathode ignition. The keeper impinges directly on the flow out of the cathode orifice and its geometry strongly influences the product between the pressure in the orifice-keeper region and the distance between cathode and keeper, $P_{ko}\cdot D_{ko}$, a key parameter for successful cathode ignition. A representative cathode equipped with a keeper is simulated using the Direct Simulation Monte Carlo method. The numerical model is first validated with computational results from the literature, and a parametric study is then conducted. Parameters include the cathode pressure-diameter in the range of 1–5 Torr-cm and the following geometric ratios: cathode orifice-to-cathode inner radii in the range of 0.1–0.7, keeper orifice-to-cathode orifice radii in the range of 1–5, and keeper distance-to-cathode-orifice diameter, in the range of 0.5–10. It is found that, if both keeper and cathode have identical orifice radii, the flow remains subsonic in the cathode orifice-to-keeper region. In most cases, however, the flow becomes underexpanded and supersonic, and the static pressure within the cathode orifice-keeper region is, on average, 4% that of the upstream pressure value. The orifice-keeper region pressure increases with either (i) a decrease in the keeper orifice diameter or (ii) an increase in the distance between the cathode orifice plate and the keeper plate, in agreement with literature experimental data. Both trends are explained through control-volume-based conservation laws. The ratio of ignition-to-nominal mass flow rates is found to be in the range of 9–120, with a most probable value of 40, in agreement with experimental literature data. This suggests that heater-less cathode ignition at a minimum DC voltage may be achieved by increasing the input mass flow rate by a factor of 40.
The views expressed in this article are those of the authors and do not reflect the official policy or position of the Department of Defense, the U.S. Government, or any agency thereof.
The following article has been submitted to the Journal of Applied Physics. After it is published, it will be found at https://pubs.aip.org/aip/jap.
Copyright (2024) Pierre-Yves C. R. Taunay, Willca Villafana, Sangeeta P. Vinoth, Igor Kaganovich, Andrei Khodak. This article is distributed under a Creative Commons Attribution-NonCommercial 4.0 International (CC BY-NC) License. https://creativecommons.org/licenses/by-nc/4.0/
I Introduction
Thermionic, orificed hollow cathodes are sources of electrons that are widely used in space electric propulsion,Goebel, Watkins, and Jameson (2007a); Mikellides and Katz (2008); Mikellides et al. (2008); Lev et al. (2019); Lev, Alon, and Appel (2019) where they deliver electrons for the main plasma discharge and to neutralize the ion exhaust beam. These cathodes consist of a tube that is capped by an orifice plate and that contains a thermionic emitter cylinder (e.g., barium-oxide). A keeper electrode, typically with a single orifice (or multiple orificesGeorgin and McDonald (2021)), faces the main cathode orifice plate downstream of the cathode tube. The keeper is used to establish the initial plasma discharge and to protect the cathode from high-energy ions that originate in the plume of the cathode and stream back towards the cathode.
A heater may be used to bring the emitter to electron emission temperatures prior to the start of the thermionic discharge. In this approach, the emitter is heated to high temperatures ($>1,200^{\circ}$), a neutral gas (e.g., xenon) is inserted in the cathode tube, and a discharge is established between the keeper and cathode. Alternatively, a DC, Paschen-like discharge may be established between the cathode and the keeper to start the cathode. Gas is inserted in the cathode tube, and a glow discharge produced by a high-voltage ($>100$ V) DC source effectively heats the cathode orifice and emitter. This approach is favored for low-to-medium amperage cathodes and circumvents the use of a heater: the heater and its associated power supply are oftentimes considered to be a single point-of-failure for space electric propulsion systems and require a dedicated power supply. The accurate placement and sizing of the keeper is critical for a reliable DC discharge: an optimal value of the pressure-distance product, $P_{ko}D_{ko}$, results in a minimum in the applied voltage required for the glow discharge, where $P_{ko}$ is the pressure of the gas in the orifice-keeper region, and $D_{ko}$ is the (axial) distance between the keeper and the cathode orifice plate. In the context of spacecraft electric propulsion, minimizing the applied voltage is desirable as it reduces the amount of on-board electric insulation required when using high-voltage power supplies.Lev et al. (2019) Although the keeper-orifice geometry is not identical to that of a Paschen discharge (e.g., the cathode orifice and keeper are not truly planar surfaces), the minimum pressure-distance product is typically $\sim$ 1–5 Torr-cm.Lev et al. (2019)
An orificed hollow cathode that operates without a plasma discharge is similar to a supersonic neutral beam source: the cathode consists of a high-pressure vessel (relative to the exhaust vessel, $P_{c}d_{c}\sim$1–10 Torr-cm, where $P_{c}$ is the cathode upstream pressure and $d_{c}$ its internal diameter) that expands into a vacuum background ($P_{0}\sim$$\mu$Torr). Because both the mean free path in the plume and the ratio of cathode-to-background pressure are large ($\geq 10^{3}$), shocks may be diffuse and the typical structure observed in supersonic neutral beams, including the “zone of silence,” may not appear.Campargue (1984) The presence of a keeper, similar to a “skimmer” in supersonic neutral beams, affects the structure of the flow downstream of the orifice.
Beyond the orifice and keeper, the neutral gas density has a direct effect on the generation of high-energy ions when the cathode operates with a plasma. Locally increasing the neutral density by injecting neutral gas in front of the keeper yields a decrease in energy of the high-energy ions generated in the plume.Chu, Goebel, and Wirz (2013); Goebel and Chu (2014) The neutral density field in this region is a critical quantity to accurately represent collisions between neutral and charged particles in numerical approaches such as particle-in-cell (PIC).
It is challenging to model the gas flow through the cathode to obtain a reliable value of $P_{ko}$ and study the flow structure beyond the orifice and keeper: the flow is compressible and rarefied, with Knudsen numbers, Kn, above $0.1$.Taunay, Wordingham, and Choueiri (2022a) For monatomic gases, the Mach number might be greater than one at the orifice exit, Lilly et al. (2004); Varoutis et al. (2008) indicating that a transition from subsonic to supersonic flow also exists within the orifice. Theoretical descriptions of the neutral flow through cathodes and/or tubes include isentropic,Domonkos (1999) Poiseuille,Goebel and Katz (2008); Becatti, Conversano, and Goebel (2021), and free-molecular flowCai and Boyd (2007) models, as well as empirical approaches.Santeler (1986) Fundamental assumptions on which the isentropic or Poiseuille flow descriptions rely are invalid in the flow regime in which cathodes operate.Taunay, Wordingham, and Choueiri (2022a) While the empirical approach suggested by SantelerSanteler (1986) provides transitional-flow corrections to the cathode pressure, it cannot describe the flow field in the orifice and in the plume. The collisionless approach of Cai and BoydCai and Boyd (2007) is valid in the plume of the cathode, where the flow is nearly free-molecular. However, it cannot describe the transitional flow in the orifice and is unable to capture the effect of the keeper on the flow past the orifice.
Cathode numerical studies that feature both charged and neutral particles include continuum approaches and/or hybrid-continuum approaches, Mikellides et al. (2005, 2006); Mikellides (2009); Mikellides et al. (2014); Sary, Garrigues, and Boeuf (2017) wherein the neutrals are in the continuum regime within the cathode tube and in the molecular flow regime in the plume, as well as Direct Boltzman solversVazsonyi, Hara, and Boyd (2020a, b); Vazsonyi (2021) and PIC approaches.Cao et al. (2019) The assumption of a continuum and free-molecular flow are valid within the dense upstream emitter region of the cathode and in the cathode plume, respectively. However, it is invalid both in the orifice and in the orifice-keeper region: the flow is transitional in both regions. Direct Boltzmann solvers, unlike PIC methods, are not subject to statistical noise. However, they are typically limited to the plume region because of their computational expense when applied to the dense and collisional internal region of the cathode. The study of Cao et al.Cao et al. (2019) features both a PIC approach for the charged particles and the Direct Simulation Monte Carlo (DSMC) technique to solve for the neutral quantities (e.g., density). However, Cao et al.Cao et al. (2019) did not include the keeper electrode as part of the simulation domain and focused only on the generation of a neutral flow field that can be used as input to a Particle-In-Cell (PIC) solver for a single test case.
In general, past DSMC studies solely dedicated to hollow cathodes have been limited. While not strictly targeted at hollow cathodes, the study of Varoutis et al.Varoutis et al. (2008) includes a wide range of orifice aspect ratio (i.e., ratio of cathode orifice length to orifice diameter) and is broadly applicable to keeper-less cathodes. Good agreement with experimental data has been reported for keeper-less cathodes.Nikrant, Jain, and Adams (2021) Daykin-Iliopoulos et al.Daykin-Iliopoulos et al. (2015) focused on the keeper-orifice region in enclosed-keeper cathodes and considered a wide range of parameters that consist of cathode-to-keeper distance, keeper-to-orifice diameter ratios, and orifice-to-insert diameter ratios. Daykin-Iliopoulos et al. observed that the pressure within the keeper-orifice region increases with increasing electrode separation and decreasing keeper diameter. However, this study did not provide a physical explanation of the observed trends.
In this work, we choose to revisit the study of Daykin-Iliopoulous et al. and explore, in depth, the effect of the keeper electrode on the fluid flow within orificed hollow cathodes that feature an enclosed keeper. Using a DSMC approach, we seek to identify which geometrical parameters most significantly impact the pressure-distance product in the orifice-keeper region of cathodes, provide broadly applicable design rules, and explain, physically, the observed trends. We first describe our numerical approach and validate our model based on reported literature data in Section II. We then perform a parameterized study based on set ranges of non-dimensional parameters that are reported in Section III. Scaled results are presented and discussed in Section IV.
II Implementation
II.1 Direct Simulation Monte Carlo
The DSMC method is a numerical technique introduced by BirdBird (1994) that may be used to simulate gases that are in the non-continuum regime. The DSMC method aims to solve numerically the Boltzmann equation with a probabilistic, particle-based approach: each numerical “particle” represents a large number of physical particles, and particle-particle collisions are handled with a probabilistic representation. The method has been applied across a variety of engineering and scientific domains, such as the aerodynamics of the Space Shuttle OrbiterRault (1994) and the fluid flow in micro and nano devices.Fang and Liou (2002)
II.2 Software
We use the DSMC code “Stochastic PArallel Rarefied-gas Time-accurate Analyzer” (SPARTA)Plimpton et al. (2019); SPA (2023) for our simulations. SPARTA is a scalable and open-source DSMC code, parallelized using Message-Passing Interface (MPI). The software has been extensively validatedGallis et al. (2015, 2016); Gu et al. (2019) and is commonly used in the DSMC community.
DSMC results are validated with a separate fluid solver, ANSYS-CFX. ANSYS-CFX is based on the continuum approach (i.e., mass, momentum, and energy equations are solved). ANSYS-CFX uses an extended upstream region which differs from that of the DSMC analysis. If both software are applied to the same region, the results would be in closer quantitative agreement. However, for simplicity and computational expediency, the DSMC code is only applied to a truncated domain. This is discussed in Section III.B.
II.3 Numerical model
II.3.1 Physical domain and boundary conditions
Figure 1 shows the physical domain representing a cathode equipped with an enclosed keeper, along with the relevant physical parameters. While some cathodesPolk et al. (1999); Becatti, Conversano, and Goebel (2021) feature a chamfered orifice, we choose not to include it for simplicity.
Wall boundary condition
Particles may interact with the walls through specular or diffuse collisions. Specular reflections result in a sign change for the component of the particle velocity that is perpendicular to the wall. The component of the particle velocity that is parallel to the wall remains unchanged, which results in all odd-order moments perpendicular to the wall vanishing. The particles cannot transfer momentum nor heat — the flow is effectively inviscid.
We choose to apply diffuse boundary conditions throughout, with an accommodation coefficient of 0.8, and do not consider the influence of gas-surface interactions on our solution. Other studiesSharipov (2004) have shown, for example, that the flow rate through an orifice is not sensitive to the value of the accommodation coefficient for vessels separated by a thin orifice.
Outflow boundary condition
Particles that reach the edge of the domain where an outflow condition is imposed simply vanish. They are not re-injected into the domain.
Inflow boundary condition
The upstream surface features a subsonic inlet with a set pressure and temperature. A varying number of particles is injected to enforce the desired pressure and temperature. The total number of injected particles accounts for domain particles that exit the simulation domain through the upstream surface.
Figure 1: Cathode computational representation.
II.3.2 Collision model
Particle-particle collisions are represented by the variable soft sphere (VSS) model. The corresponding VSS parameters are shown in Table 1 for xenon.
Table 1: VSS parameters for xenon (Ref. Bird, 1994, pp. 408,410)
Species $d_{\textrm{ref}}$(m) $\omega$ $T_{\textrm{ref}}$ $\alpha$
Xenon $5.65\times 10^{-10}$ 0.85 273.15 1.44
II.3.3 Discretization
The time step, $\Delta t$, cell size, $\Delta x$ and $\Delta y$, and particle weight (i.e., the ratio of real to simulated particles), $f_{n}$, must be fine enough to accurately represent collisional processes in a DSMC calculation. The collision mean free path, $\lambda$, collision time scale, $t_{c}$, and the particle transit time, $t_{p}$, shall all be accurately resolved, even for cells with the highest densities and particle velocities:
  • $\Delta x\approx\Delta y\approx\lambda/N_{\lambda}$, and
  • $\Delta t<\textrm{min}\left(t_{c}/N_{t},t_{p}/N_{t}\right)$,
where $2\leq N_{\lambda},N_{t}\leq 3$. The particle weight shall be chosen such that there is a statistically significant number of particles in most simulation cells, even for those with the lowest density. In practice, this means that there should be at least 20 particles per cell. The total number of particles per cell is verified after a given simulation is complete. Figure 2 shows an example distribution of the number of particles per cell for a simulation with a locally refined grid. While it is clearly multimodal due to the large variations in density between the cathode channel and the plume, the number of particles per cell remains statistically significant in most cells. For this particular example, the 5% and 95% bounds of the distribution are 53 and 297 particles per cell, respectively, with the most probable peaks located at 22, 186, and 296 particles per cell.
Figure 2: Illustrative distribution and whisker plot of the total number of simulated particles per cell.
Space discretization
The number of cells in the $j$-th direction is
$N_{j}\approx\dfrac{L_{j}}{\lambda/N_{\lambda}},$ (1)
where $L_{j}$ is the domain length in that direction and $\lambda$ is the collision mean free path of neutral atoms. The total number of cells in a given direction is rounded to the nearest power of two. The mean free path for the VSS model, $\lambda$, is given by
$\lambda=\dfrac{1}{\sqrt{2}n_{n}\sigma_{T}},$ (2)
where $n_{n}$ is the local neutral density and $\sigma_{T}$ is the total collision cross section. The cross section for the VSS model is given byBird (1994) (pp. 40–42):
$\sigma_{T}=\pi d_{\textrm{ref}}^{2}\left(\dfrac{T_{\textrm{ref}}}{T}\right)^{\omega-1/2},$ (3)
where $T$ is the local temperature.
Time step
The inter-particle collision and transit time scales are given by
$t_{c}=\dfrac{1}{n_{n}\sigma\bar{v}},\textrm{and}$ (4)
$t_{p}=\dfrac{\textrm{min}\left(\Delta x,\Delta y\right)}{\bar{v}},$ (5)
where $\bar{v}$ is the most probable velocity. Assuming that particles follow a Maxwellian distribution,
$\bar{v}=\sqrt{\dfrac{2k_{B}T_{n}}{M}},$ (6)
[…]
Particle weight
In axisymmetric coordinates, the volume of a given cell is given by
$V_{j}=\pi\left((j\Delta y)^{2}-(j-1)\Delta y)^{2}\right)\Delta x,$ (7)
where $j$ is the index of the cell along the $y$-axis. The corresponding number of real and simulated particles is $V_{c}n_{n}$ and ${V_{c}n_{n}/f_{n}}$ respectively. The particle weight is chosen such that the cell with the smallest volume (i.e., the center cell in axisymmetric coordinates) has at least 20 simulated particles, even in the lowest-density regions. Because the volume of the axisymmetric cell increases with radius, cells on the outermost boundary of the domain will contain a larger number of simulated particles. To ensure that each cell contains an adequately balanced number of particles, we assign a radius-based weight to the cells, which internally rescales $f_{n}$ by the local cell volume in SPARTA.
II.3.4 Convergence and statistics
The domain is initialized with a non-zero number density that is computed based on a vacuum condition of 1 $\mu$Torr and a gas temperature of 300 K (i.e., $n_{n}=3.2\times 10^{16}$ m-3). Particles are drawn from a Maxwellian distribution with zero streaming velocity and the upstream temperature.
The simulation is considered to have converged once the total number of particles remains approximately constant. Figure 3 shows the typical evolution of the total number of (simulated) particles as a function of time. Because the initial density in the cathode is much lower than the steady-state one (approximately two-to-three orders of magnitude) the total number of particles increases until it reaches steady-state.
Figure 3: Typical evolution of the total number of particles.
Relevant physical quantities are computed once steady-state is reached. Results are averaged over 10,000 time steps ($\sim$ 0.1 ms) to decrease the statistical noise inherent to particle simulations.
II.4 Multi-level grid and load rebalance
The time-to-solution is reduced by (i) decreasing the overall number of cells within the domain with a multi-level grid and (ii) rebalancing the workload across multiple processors as particles appear within or vanish from the domain. The grid is refined within the cathode tube and keeper regions. A coarse grid is sufficient for the plume region. An illustrative, adapted grid is shown in Figure 4. Care is taken to not extensively refine the low-density region past the corner of the keeper exit. Because the particle weight, $f_{n}$, is set globally for the simulation, a refined grid in this region would result in a low number of DSMC particles per computational cell.
Figure 4: Illustrative adapted grid for the case where $d_{k}/d_{o}=1$ and $D_{ko}/d_{o}=0.5$. The cathode surface is indicated in gray. The internal grid is 32$\times$ more refined than the plume grid.
Because the number of DSMC particles increases significantly within the first 10,000 steps of the simulation (see, e.g., Figure 3), we rebalance the particle load across processors every 100 steps during that time period. Periodic rebalancing is applied throughout the simulation thereafter.
II.5 Numerical validation
We validate the DSMC model with the test case described by Cao et al.,Cao et al. (2019) which features a computational domain without a keeper electrode. A fully coupled PIC/DSMC approach wherein both charged and neutral particles are simulated is beyond the scope of this article: we cannot capture plasma-neutral effects such as the increase in neutral particles temperature near the orifice due to the frequent charge-exchange collisions with the plasma ions.
The parameters describing the computational domain of the validation case are shown in Table 2. Using the perfect gas law and the provided upstream pressure and density (450 Pa and 1.3$\times$1022 m-3, respectively), the temperature of the injected gas is estimated to be 2,500 K. Unlike Cao et al., we do not rescale the heavy particle mass or domain size. For the set input upstream conditions, this results in an overestimation of the mass flow rate within the cathode. However, quantities such as the (static) pressure and the number density can be directly compared.
The data reported by Cao et al.Cao et al. (2019) indicate that the neutral gas temperature is kept constant throughout the domain. We apply a wall temperature equal to that of the injected particles (i.e., $\approx$2,500 K) to replicate Cao et al.’s results within the internal cathode region. However, there is no mechanism by which we can enforce a constant static temperature in the entire domain.
Table 2: Physical domain for the test case presented in Cao et al.Cao et al. (2019)
Parameter Unit Symbol Value
Inner channel diameter mm $d_{c}$ 5
Orifice diameter mm $d_{o}$ 2
Orifice length mm $L_{o}$ 1
Domain length mm $L_{d}$ 16
Domain radius mm $r_{d}$ 4
Upstream pressure Pa $P_{n}$ 450
Upstream density m-3 $n_{n}$ $1.3\times 10^{22}$
Upstream temperature K $T_{n}$ 2,500
Neutral species mass amu $M_{a}$ 131.293
Under the assumption that the flow is choked and reaches Mach 1 at the orifice exit, the static pressure at the exit plane of the orifice, $P_{\textrm{exit}}$, may be estimated with either an isentropic flow or a Fanno flow model. In both cases, the static pressure is given by
$P_{\textrm{exit}}=P_{\textrm{T,exit}}\left(\dfrac{2}{\gamma+1}\right)^{\gamma/\left(\gamma-1\right)},$ (8)
[…]
$\dfrac{P_{\textrm{T,exit}}}{P_{\textrm{T,in}}}=\dfrac{1}{M_{in}}\left[\dfrac{2+\left(\gamma-1\right)M_{in}^{2}}{\gamma+1}\right]^{\dfrac{\gamma+1}{2\left(\gamma-1\right)}},$ (9)
where the subscripts ${in}$ and ${exit}$ denote the orifice entrance and exit quantities, respectively. For this particular validation case, the entrance Mach number is found to be $M_{in}\approx 0.5$ from the computational results. Under the assumption that the flow is isentropic in the cathode emitter region, $P_{\textrm{T,in}}=P_{\textrm{T}}$, and Equation 8 may then be used to obtain $P_{\textrm{exit}}/P_{T}$.
A Poiseuille flow approach is often used to estimate the pressure within the cathode.Goebel and Katz (2008); Mizrahi et al. (2012); Becatti, Conversano, and Goebel (2021); Becatti and Goebel (2022); Poli et al. (2023) We use the model from Goebel and KatzGoebel and Katz (2008) with the xenon viscosity from Stiel and ThodosStiel and Thodos (1961). The mass flow rate may be estimated by assuming a continuum flow and a choked orifice,
$\dot{m}=C_{d}\dfrac{P_{T}\pi r_{o}^{2}}{\sqrt{T_{T}}}\sqrt{\dfrac{\gamma}{R_{g}}}\left(\dfrac{2}{\gamma+1}\right)^{\left(\gamma+1\right)/\left(\gamma-1\right)}$ (10)
where $C_{d}$ is the discharge coefficient and $R_{g}$ the gas constant.
Figure 5 compares the pressure and the neutral density as obtained with our theoretical and numerical approaches to that of Cao et al. We also compare our results to the collisionless free-jet solution obtained by Cai and BoydCai and Boyd (2007) and to the results of the fluid solver ANSYS-CFX. The average orifice exit (static) temperature, gas velocity, and density scale required for the analytical model of Cai and BoydCai and Boyd (2007) are derived from our simulation. They are equal to 1,823 K, 447 m/s, and $5.0\times 10^{21}$ m-3, respectively. Reasonable agreement is obtained for all quantities for all numerical approaches. Because the flow is transitional in the orifice and near-orifice region (i.e., Knsimilar-$\sim$0.1–1), our numerical results differ from both the collisionless solution and the fluid solution. The DSMC and collisionless approaches yield the same density decay in the far field ($z\geq 12$ mm) where the Knudsen number increases. However, the fluid solver is unable to resolve this density decay. The Fanno flow model of the orifice yields the most accurate theoretical value. The Poiseuille flow approach overestimates the pressure and is, in general, inaccurate. As discussed elsewhere,Taunay, Wordingham, and Choueiri (2022a) this approach should not be used to estimate the pressure within cathodes as it is inapplicable to the flow regime in which cathodes operate: most of the assumptions required by this flow model are invalid in cathodes. For example, the flow is not fully developed in the orifice and compressibility effects are significant since the flow becomes sonic at the orifice exit.
Figure 5: On-axis (a) neutral pressure and (b) density for Cao et al.’s test case. All flow models start with the same value at $z=0$ mm. The isentropic and Fanno flow models are assumed to have the same static quantities at the orifice entrance. The error bars on the Poiseuille flow model indicate values of $C_{d}$ in the range of 0.1–1, with the symbol set at the pressure and density values for which $C_{d}=0.5$.
Figure 6(a) shows the neutral density contour for this test case. The Knudsen number (defined here with the characteristic length equal to the orifice diameter, $L=d_{o}$) is shown in Figure 6(b) and illustrates that, for this particular test case, the cathode operates in the transitional ($\textrm{Kn}\sim 0.1$) to free molecular ($\textrm{Kn}\geq 10$) regimes.
[…]
Figure 6: Contours of (a) neutral density and (b) Knudsen number (normalized by the orifice diameter) for Cao et al.’s test case.
A keeper electrode located at $\sim 1$ mm downstream of the orifice would impinge on the flow and affect local, static quantities. We explore the effect of a keeper electrode placement and size in the next sections.
III Parametric study
III.1 Scope
We now consider a representative hollow cathode with an inner diameter of 0.5 cm and investigate the influence of the keeper electrode on the neutral flow. Most cathodes operate in a limited range of internal pressure-inner diameter products of 1–10 Torr-cm.Taunay, Wordingham, and Choueiri (2022b) During operation, the neutral gas is heated by frequent charge-exchange collisions and reaches temperatures that are estimated to be 2,000–4,000 K.Goebel and Katz (2008) Because this work is concerned with the neutral flow prior to ignition, the upstream temperature of the neutral gas is, instead, set to a value of $T_{T}=300$ K. The neutral density within the cathode and orifice is, therefore, higher than that of the high-temperature cases, and the Knudsen number is, correspondingly, lower. The Knudsen number in the cathode may be evaluated by assuming that the cathode orifice is choked, with a Mach number of $1$ within the orifice. The corresponding static temperature in the orifice is 225 K. Figure 7 shows the corresponding range of insert and orifice Knudsen numbers for this representative cathode, computed with Equation 2 for the mean free path and using the cathode insert or orifice diameter as the relevant length scale. In most cases, the flow can be assumed to be continuum within the insert and orifice region.
Figure 7: Range of insert and orifice Knudsen numbers for the representative cathode.
Because we are considering only the neutral gas with an upstream temperature of $T_{T}=300$ K, we limit our study to pressure-diameters of 1–5 Torr-cm. For the cathode geometry considered, this corresponds to an upstream pressure, $P_{T}$, of 2–10 Torr (266–1330 Pa).
We further limit our study to enclosed-keeper cathodes: the boundary between the orifice and the keeper is considered to be a wall. All other boundary conditions remain the same as described in Section II.
All dimensions may be rescaled by the inner cathode diameter, $d_{c}$, and the orifice diameter, $d_{o}$. The range of parameters studied was determined based on the dimensions of cathodes that are reported in the literature. Figure 8 and Table 3 show the range of rescaled parameters investigated.
Figure 8: Range of rescaled parameters.
The cathodes for which dimensions are available, along with their respective dimensions, are shown in Table 4, in the Appendix.
Table 3: Range of scaled parameters considered
Parameter Unit Symbol Value
Inner channel diameter mm $d_{c}$ 5
Orifice diameter $d_{o}/d_{c}$ 0.1–0.7
Orifice length $L_{o}/d_{o}$ 1
Keeper distance $D_{ko}/d_{o}$ 0.5, 1–3, 10
Keeper diameter $d_{k}/d_{o}$ 1–5
Keeper length $L_{k}/d_{o}$ 1
Pressure-diameter product Torr-cm $P_{T}d_{c}$ 1, 5
Upstream pressure Pa $P_{T}$ 266, 1330
Upstream temperature K $T_{T}$ 300
Neutral species mass amu $M_{a}$ 131.293
Using Equation 10 and assuming $C_{d}=1$, the range of mass flow rates considered is approximately 0.28–13.5 mg/s (2.8–135 sccm of xenon). We limit our study to $d_{o}/d_{c}=$ 0.1 for $P_{T}d_{c}=5$ Torr-cm: for this particular pressure-diameter, the corresponding mass flow rate is above 34 mg/s (i.e., $\geq 300$ sccm) for $d_{o}/d_{c}=0.5$, which is much higher than typical cathode flow rates.
III.2 Computational domain
In order to decrease the overall computational cost and improve the time-to-solution we restrict the size of the domain, adapt the grid, and consider that the size of single computational cell is half that of the smallest mean free path in the densest regions of the simulation domain (i.e., $N_{\lambda}=2$).
For most cases, we do not consider the high-density insert region and introduce particles through the orifice exit directly. However, care must be taken in specifying the profile of the flow quantities at the orifice exit. For a temperature of 300 K, the Reynolds number at the orifice exit, $\textrm{Re}_{o}$, is on the order of $\mathcal{O}\left(10^{2}\right)$. While this places the flow in the laminar regime, it is not fully developed at the orifice exit. The entrance length, $L_{e}$, or length before which the flow is fully developed, is given by:White (2015)
$\dfrac{L_{e}}{d_{o}}=\dfrac{0.6}{1+0.035\textrm{Re}_{o}}+0.0575\textrm{Re}_{o}.$ (11)
In all cases, $\dfrac{L_{e}}{d_{o}}>\dfrac{L_{o}}{d_{o}}$: the flow is not fully developed.
The Mach number is typically in the range of 0.9–1.2 at the orifice exitLilly et al. (2004); Varoutis et al. (2008); Titarev and Shakhov (2014) due to slip flow, viscosity, and 2-D effects. We have observed a Mach number at the orifice inlet, $M_{o,in}$, of $\approx 0.4$ in our simulations. To determine the orifice exit conditions, we assume that the flow
  • may be modeled with a Fanno flow approach (i.e., the flow is in the continuum regime in the orifice), and
  • is sonic at the orifice exit (Ma$=1$).
The Fanno flow friction coefficient, $f$, for the orifice is given by:Anderson (2003)
$f=\dfrac{1}{4}\dfrac{d_{o}}{L_{o}}\left(\dfrac{1-M_{o,in}^{2}}{\gamma M_{o,in}^{2}}+\dfrac{\gamma+1}{2\gamma}\log\left(\dfrac{\left(\gamma+1\right)M_{o,in}^{2}}{2+\left(\gamma-1\right)M_{o,in}^{2}}\right)\right).$ (12)
For an aspect ratio $L_{o}/d_{o}=1$ and $M_{o,in}\approx 0.4$, $f\approx 0.5$.
A “flat-top” velocity, temperature, and density profile is typically not observed in orifice flows in the hydrodynamic regime. Because the flow is not fully developed, the radial velocity profile, therefore, does not follow the usual square dependency on the radius that laminar flows exhibit. Figure 9 shows the radial profiles of axial velocity, temperature, and density, for the case where $P_{T}d_{c}=1$ Torr-cm, $d_{o}/d_{c}=0.1$, $D_{ko}/d_{o}=1$, and $d_{k}/d_{o}=2$.
Figure 9: Normalized radial profile of (a) axial velocity, (b) Mach number, (c) number density, and (d) static temperature, for the case where $P_{T}d_{c}=1$ Torr-cm, $d_{o}/d_{c}=0.1$, $D_{ko}/d_{o}=1$, and $d_{k}/d_{o}=2$.
Velocity profiles typically used for turbulent flows fit the DSMC results reasonably well: both the “1/7-th” power law and the sinusoidal fit suggested by De Chant De Chant (2005) show good agreement with our numerical data. Figure 10 shows the center-line Mach number, pressure, density, and axial velocity, for the case where $P_{T}d_{c}=1$ Torr-cm, $d_{o}/d_{c}=0.1$, $D_{ko}/d_{o}=1$, and $d_{k}/d_{o}=2$. Three cases are shown:
  • a complete solution that includes the high-density insert region,
  • a solution for a truncated domain with a flat-top profile at the cathode orifice exit (i.e., the inlet for this particular case), and
  • a solution for a truncated domain with an imposed, non-flat velocity and temperature profile at the cathode orifice exit.
The numerical results for the simplified domain are qualitatively similar to that of the complete domain. However, the Mach number, pressure, and velocity are all under-estimated near the cathode orifice exit ($x/d_{o}<1$) for the truncated domain cases. This is likely a result of (i) the radial velocity being neglected near the orifice wall and (ii) the orifice exit density being underestimated with the Fanno flow approach. However, good agreement is obtained with the full solution for $x/d_{o}>0.5$ when the inlet velocity follows a sinusoidal radial profile.
Figure 10: Center-line (a) axial velocity, (b) Mach number, (c) number density, and (d) static temperature for the case where $P_{T}d_{c}=1$ Torr-cm, $d_{o}/d_{c}=0.1$, $D_{ko}/d_{o}=1$, and $d_{k}/d_{o}=2$. Numerical results for both a complete domain and a truncated domain are shown. A flat-top profile for the velocity, density, and temperature is assumed for the inlet of the truncated domain.
For the truncated domain cases, particles will, therefore, be introduced in the keeper region directly from the orifice. The velocity profile will be assumed to be sinusoidal. Assuming a friction coefficient, $f$, equal to 0.5, and an orifice exit Mach number of one, the inlet flow quantities are determined from the upstream stagnation quantities through the following procedure:
  1. 1.
    Using a Fanno flow approach, solve for Mach number and relevant pressure and temperature ratios at the orifice entrance.
  2. 2.
    Compute the pressure, temperature at the orifice exit from the ratios obtained in Step 1 and the assumed Mach number of 1 at the orifice exit
  3. 3.
    Compute the radial velocity profile using the approach suggested by DeChantDe Chant (2005)
  4. 4.
    Compute the radial static temperature profile from the velocity profile and the definition of the Mach number.
IV Results and discussion
IV.1 Flow structure in the presence of a keeper electrode
Accurately resolving the neutral flow field is critical for comprehensive plasma-neutral cathode simulations that use the DSMC method for the neutrals. The introduction of an enclosed keeper strongly affects the neutral flow downstream of the orifice. The flow structure differs from that of a simple expansion from the cathode orifice into vacuum (see, e.g., Figure 6(a)). In the presence of the keeper, the flow through the orifice first expands into a larger channel (i.e., the orifice-keeper region). After reaching the keeper, the flow undergoes a second expansion, into vacuum. Our numerical results reveal the existence of two main flow regimes: the flow may be either supersonic or subsonic in the orifice-keeper region.
IV.1.1 Supersonic flow
In most cases, the pressure ratio $P_{ko}/P_{T}$ is such that the flow that leaves the cathode orifice is choked and supersonic, and the overall flow structure is similar to that reported by Titarev and ShakovTitarev and Shakhov (2014) for a composite pipe system composed of two pipes of increasing radii that exhaust into vacuum. Figures 11 through 14 illustrate the typical flow structure in the case where the orifice is choked. The flow expansion from the cathode orifice to the keeper-orifice region results in an underexpanded, rarefied, supersonic jet (Figures 111214). After exiting the orifice, the jet interacts with the keeper plate and a toroidal zone of recirculation appears in the region between the underexpanded jet and the top boundary (Figure 11). The Knudsen number in the keeper region is $\approx 0.1$ (Figure 13), indicating that the flow is in the transitional regime.
Refer to captionOrifice
Figure 11: Velocity magnitude with streamlines for the case $P_{T}d_{c}=1$ Torr-cm, $d_{o}/d_{c}=0.3$, $D_{ko}/d_{o}=3.0$, and $d_{k}/d_{o}=2.0$.
Refer to captionOrifice
Figure 12: Mach number for the case $P_{T}d_{c}=1$ Torr-cm, $d_{o}/d_{c}=0.3$, $D_{ko}/d_{o}=3.0$, and $d_{k}/d_{o}=2.0$.
[…]
Figure 13: Knudsen number normalized by orifice diameter for the case $P_{T}d_{c}=1$ Torr-cm, $d_{o}/d_{c}=0.3$, $D_{ko}/d_{o}=3.0$, and $d_{k}/d_{o}=2.0$.
Refer to captionOrifice
Figure 14: Static pressure for the case $P_{T}d_{c}=1$ Torr-cm, $d_{o}/d_{c}=0.3$, $D_{ko}/d_{o}=3.0$, and $d_{k}/d_{o}=2.0$.
Figure 15: Velocity magnitude with streamlines for the case $P_{T}d_{c}=1$ Torr-cm, $d_{o}/d_{c}=0.3$, $D_{ko}/d_{o}=10.0$, and $d_{k}/d_{o}=2.0$.
The exact structure of the underexpanded jet depends on the backing-to-upstream pressure ratio, and, therefore, the cathode geometry. The centerline Mach number, static pressure, and number density are shown in Figures 16 through 19 for $d_{o}/d_{c}=$ 0.1–0.7, respectively. We observe the following qualitative trends:
  • For short orifice-to-keeper distances and larger keeper-to-orifice ratios ($D_{ko}/d_{o}\leq 1$, $d_{k}/d_{o}\geq 3$), the orifice and keeper expansions merge to form a single expansion system (see, e.g., Figure 17(d) for $D_{ko}/d_{o}=0.5$).
  • For the longest orifice-to-keeper distances ($D_{ko}/d_{o}=10$), the orifice expansion results in a subsonic flow in the orifice-keeper region, and the recirculation zone remains confined near the cathode orifice plate (Figure 15). The flow transitions from supersonic to subsonic in the orifice-keeper region and a plateau in both density and pressure exist. The transition may be through an expansion ($d_{o}/d_{c}=0.1$) or through a diffuse shock ($d_{o}/d_{c}\geq 0.3$). In the latter case, the static pressure increases past the shock (see, e.g., Figure 18(a–c) for $x/D_{ko}\approx 0.3$).
  • As the keeper-to-orifice diameter ratio, $d_{k}/d_{o}$, increases, the keeper expansion into vacuum becomes more pronounced: at the keeper entrance, the Mach number increases with $d_{k}/d_{o}$, while both static pressure and number density decrease with $d_{k}/d_{o}$ (see, e.g., Figures 19(a,d,g))
  • For large enough mass flow rates (i.e., large enough orifice-to-cathode diameter ratio since $P_{T}$ is fixed), multiple cells of the underexpanded jet appear and a Mach disk may form (see, e.g., Figure 19(c), for $D_{ko}/d_{o}=3$).
Figure 17(a–c) show a comparison of the DSMC numerical results to that produced by ANSYS-CFX, for select cases ($d_{o}/d_{c}=0.3$, $d_{k}/d_{o}=2$, and $D_{ko}/d_{o}=0.5,\,3,\,\textrm{and}\,10$). Results from the continuum flow solver qualitatively reproduce that of the DSMC solver. However, the values for the density (and pressure) are consistently underestimated (resp., overestimated), and both density and pressure decay rates in the keeper and plume regions are inconsistent with that predicted by the DSMC approach. The continuum flow solver solution was obtained on a computational domain that extends upstream in the insert region, which was not covered by the DSMC analysis. Therefore, the inlet parameters of the DSMC analysis presented on Figure 9 differ from the profiles obtained with a consistent analysis that uses a continuum model. Better closer quantitative agreement between the DSMC and continuum models may be obtained if the DSMC analysis is performed using initial profiles at the nozzle that correspond to the results of the continuum flow analysis. This approach is beyond the scope of this work.
IV.1.2 Subsonic flow
In some cases where $d_{o}=d_{k}$, the keeper induces a (static) pressure rise that is large enough to prevent the orifice from choking. In this configuration the flow remains subsonic throughout the orifice-keeper region, and the supersonic expansion only occurs in the keeper orifice (as opposed to the main cathode orifice). Because the flow is subsonic in this particular case, the keeper-to-total pressure ratio, $P_{ko}/P_{T}$, is higher than that in the supersonic case. Figures 20 and 21 show a contour of the Mach number and static pressure for the subsonic case, respectively. For high-enough mass flow rates, however, the flow may be supersonic. The exact point at which the flow transitions from subsonic to supersonic for the case where $d_{o}=d_{k}$ is beyond the scope of this article.
Figure 16: Centerline values for $d_{o}/d_{c}=0.1$ and $P_{T}d_{c}=1$ Torr-cm. (a,d,g) Normalized density, (b,e,h) normalized pressure, (c,f,i) Mach number. Density and pressure are normalized by the upstream values ($n_{T}=6.4\cdot 10^{22}$ m-3 and $P_{T}=267$ Pa, respectively). The horizontal axis is normalized by the orifice-keeper distance, $D_{ko}$.
Figure 17: Centerline values for $d_{o}/d_{c}=0.3$ and $P_{T}d_{c}=1$ Torr-cm. (a,d,g) Normalized density, (b,e,h) normalized pressure, (c,f,i) Mach number. Density and pressure are normalized by the upstream values ($n_{T}=6.4\cdot 10^{22}$ m-3 and $P_{T}=267$ Pa, respectively). The horizontal axis is normalized by the orifice-keeper distance, $D_{ko}$.
Figure 18: Centerline values for $d_{o}/d_{c}=0.5$ and $P_{T}d_{c}=1$ Torr-cm. (a,d,g) Normalized density, (b,e,h) normalized pressure, (c,f,i) Mach number. Density and pressure are normalized by the upstream values ($n_{T}=6.4\cdot 10^{22}$ m-3 and $P_{T}=267$ Pa, respectively). The horizontal axis is normalized by the orifice-keeper distance, $D_{ko}$.
Figure 19: Centerline values for $d_{o}/d_{c}=0.7$ and $P_{T}d_{c}=1$ Torr-cm. (a,d,g) Normalized density, (b,e,h) normalized pressure, (c,f,i) Mach number. Density and pressure are normalized by the upstream values ($n_{T}=6.4\cdot 10^{22}$ m-3 and $P_{T}=267$ Pa, respectively). The horizontal axis is normalized by the orifice-keeper distance, $D_{ko}$.
Figure 20: Mach number for the (subsonic) case $P_{T}d_{c}=1$ Torr-cm, $d_{o}/d_{c}=0.3$, $D_{ko}/d_{o}=3.0$, and $d_{k}/d_{o}=1.0$.
Figure 21: Static pressure for the (subsonic) case $P_{T}d_{c}=1$ Torr-cm, $d_{o}/d_{c}=0.3$, $D_{ko}/d_{o}=3.0$, and $d_{k}/d_{o}=1.0$.
IV.2 Keeper pressure-distance product
We now turn our attention to the pressure-distance product in the orifice-keeper region, a critical quantity for successful cathode ignition. The pressure-distance product in the keeper-orifice region is given by
$P_{ko}D_{ko}=\dfrac{P_{ko}}{P_{T}}\dfrac{D_{ko}}{d_{o}}\dfrac{d_{o}}{d_{c}}P_{T}d_{c}=\beta\left(\dfrac{d_{k}}{d_{o}},\dfrac{D_{ko}}{d_{o}},\dfrac{d_{o}}{d_{c}}\right)P_{T}d_{c},$ (13)
where $P_{ko}$ is the average pressure in this region, and $\beta$ includes both the linear and non-linear effects of the geometry and mass flow rate on the pressure-diameter product with geometry. Linear factors and non-linear factors include $D_{ko}/d_{o}$, $d_{o}/d_{c}$, and $P_{T}d_{c}$; and $d_{k}/d_{o}$ (through $P_{ko}/P_{T}$), respectively.
We compute the keeper-orifice region pressure, $P_{ko}$, by averaging the static pressure of all cells (i) that are bounded by the orifice and keeper entrance planes and (ii) for which the flow remains subsonic. Those cells are part of the recirculation zone. The range of scaled average keeper-to-upstream pressure ratios, $P_{ko}/P_{T}$, is shown in Figure 22 for the entire dataset of 87 numerical cases.
Figure 22: Combined violin and box plot of the range of keeper-to-upstream pressure ratios, $P_{ko}/P_{T}$. Each individual tick mark indicates the location of a data point.
The 95% bounds of the data are 0.004 and 0.51, respectively, with a mean of 0.04. This wide range of values is due to the variety of flow conditions encountered: the highest pressure ratios ($\geq 0.35$) correspond to an entirely subsonic flow throughout the keeper-orifice region (i.e., $d_{k}/d_{o}=1$ and $P_{T}d_{c}=1$ Torr-cm), while the lowest ones ($\leq 0.01$) correspond to a strongly underexpanded jet.
Which geometrical parameters most influence the ratio of keeper-to-upstream pressures, $P_{ko}/P_{T}$? The ratio is shown in Figure 23 for the entire dataset, as a function of the keeper-to-orifice diameter ratio, $d_{k}/d_{o}$.
Figure 23: Keeper to upstream pressure for the entire dataset as a function of the keeper-to-orifice diameter ratio, $d_{k}/d_{o}$. $P_{T}d_{c}$ is set to (a) 1 Torr-cm and (b) 5 Torr-cm.
Two branches exist for $d_{k}/d_{o}=1$:
  • For $P_{T}d_{c}=1$ Torr-cm, the flow remains subsonic in the entire orifice-keeper region and the keeper-to-upstream pressure ratio, is correspondingly, close to unity.
  • For $P_{T}d_{c}=5$ Torr-cm, the flow becomes supersonic and the pressure ratio subsequently decreases.
The keeper-to-upstream pressure ratio, $P_{ko}/P_{T}$, varies inversely with $d_{k}/d_{o}$:
$P_{ko}/P_{T}\propto\left(d_{o}/d_{k}\right)^{r},$ (14)
where $r=2.55$ and $r=1.75$ for $P_{T}d_{c}=$ 1 and 5 Torr-cm, respectively. A mass, momentum, and energy balance on the control volume applied to the keeper-orifice region shows that, under the simplifying assumption of continuum flow,
$\dfrac{P_{ko}}{P_{T}}=\dfrac{1}{\left(d_{k}/d_{o}\right)^{2}-1}\left[\dfrac{P_{o}}{P_{T}}\mathcal{F}\left(M_{o},M_{k}\right)+8\dfrac{\left<\tau\right>}{P_{T}}\dfrac{d_{k}}{d_{o}}\dfrac{D_{ko}}{d_{o}}\right],$ (15)
where $M_{o}$ and $M_{k}$ are the Mach number at the orifice exit and keeper entrance, respectively, and $\tau$ is the axial wall shear stress. The function $\mathcal{F}$ is given by
$\begin{split}\mathcal{F}=&\dfrac{M_{o}}{M_{k}}\left(\dfrac{1+\dfrac{\gamma-1}{2}M_{o}^{2}}{1+\dfrac{\gamma-1}{2}M_{k}^{2}}\right)^{1/2}\cdot\left[1+\gamma M_{k}^{2}\right]\dotsm\\ &\dotsm-\left(1+\gamma M_{o}^{2}\right)\end{split}$ (16)
Equation 15 may be evaluated provided that the pressure ratio $P_{o}/P_{T}$ and the Mach number at the keeper entrance, $M_{k}$, are known. Because we assumed a Fanno flow in the orifice with an entrance Mach number of 0.4 and an orifice aspect ratio of 1, $P_{o}/P_{T}\approx 1/1.59$. Figure 24 shows the range of numerical values of $M_{k}$ obtained, as a function of the keeper-to-orifice diameter ratio. The corresponding values of $P_{ko}/P_{T}$ are shown in Figure 23 for the inviscid case (i.e., $\tau\right>=0$, for simplicity). Reasonable agreement is obtained with the numerical results, which indicates that the variation of the pressure ratio with $d_{k}/d_{o}$ can be explained through control-volume-based conservation laws.
Figure 24: Keeper entrance Mach number, $M_{k}$, as a function of the keeper-to-orifice diameter ratio, $d_{k}/d_{o}$. The dashed line indicates the average value.
Figure 25: Combined violin and box plot of the range of values of (a) $\beta$ and (b) $1/\beta$, where $\beta=\dfrac{P_{ko}}{P_{T}}\dfrac{D_{ko}}{d_{o}}\dfrac{d_{o}}{d_{c}}$, for the entire dataset (including cases for which the cathode orifice is not choked). Each individual tick mark indicates the location of a data point. The data range for (b) is limited to the interval that contains 95% of all values.
The distribution of the values of $\beta$ and 1/$\beta$ that we obtained for our parametric study is shown in Figure 25. The 95% interval of $\beta$ is $7\times 10^{-4}$$3.6\times 10^{-1}$, with an average of 0.07 and a most probable value of 0.024: the pressure-distance product in the keeper region is equal to 2.4% of the pressure-diameter product imposed in the upstream cathode region.
For a choked flow, the ratio between the ignition and nominal mass flow rates, $\dot{m}_{i}$ and $\dot{m}_{n}$, is
$\dfrac{\dot{m}_{i}}{\dot{m}_{n}}=\dfrac{P_{T,i}}{P_{T,n}}\left(\dfrac{T_{T,n}}{T_{T,i}}\right)^{1/2},$ (17)
where $P_{T}$ and $T_{T}$ are the total, upstream conditions. The subscripts $i$ and $n$ denote the ignition and nominal conditions, respectively. Assuming that the flow temperature remains constant ($T_{T,n}=T_{T,i}$), we have, using Equation 13:
$\dfrac{\dot{m}_{i}}{\dot{m}_{n}}=\dfrac{P_{T,i}d_{c}}{P_{T,n}d_{c}}=\frac{1}{\beta}\dfrac{P_{ko,i}D_{ko}}{P_{T,n}d_{c}}$ (18)
Because most cathodes operate near $P_{T,n}d_{c}\sim 1$ Torr-cmTaunay, Wordingham, and Choueiri (2022a) and with noble gases for which the ignition pressure-distance product is near the Paschen minimum (i.e., $P_{ko,i}D_{ko}\sim 1$ Torr-cm), the last term simplifies and the ratio of ignition to nominal mass flow rates is equal to the ratio between the upstream pressure-diameter product and the pressure-distance product:
$\dfrac{\dot{m}_{i}}{\dot{m}_{n}}\approx\dfrac{1}{\beta}$ (19)
If $d_{k}/d_{o}>1$, this ratio has a median of 63.5 and a median absolute deviation of 54.2. The most-probable value is $1/\beta\approx 50$. Because the flow is choked, this suggests that a fifty-fold increase in mass flow rate results in a favorable pressure-distance product for a Paschen-like discharge. Increasing the mass flow rate prior to cathode ignition is a strategy that is typically used in orificed-emitter hollow cathodes.Becatti, Conversano, and Goebel (2021); Ham et al. (2019) We show in Figure 26 the ratio of experimentally observed ignition flow rate to nominal flow rate for a few cathodes operating on both argon and xenon, along with the range of favorable flow rate ratios as determined from the median, $M$, median absolute deviation, MAD, and most-probable value.
Figure 26: Ratio of ignition to nominal mass flow rates for multiple cathodes. The median, MAD, and most probable value of $1/\beta$ are limited to the dataset for which $d_{k}/d_{o}>1$ (i.e., the cases for which the flow is choked). The experimental data were retrieved from Refs Payman and Goebel, 2022; Becatti, Conversano, and Goebel, 2021 for the HHC developed by the Jet Propulsion Laboratory (JPL) and from Daykin-Iliopoulos et al., 2015,Ning et al., 2019, and Ham et al., 2019 for those developed by the University of Southampton, Harbin Institute of Technology (HIT), Colorado State University (CSU), respectively. All cathodes operate on xenon, except for the Southampton HHC, which operates on argon.
V Conclusion
We have performed a parametric study to quantify the effect of the changes in keeper geometrical parameters on the neutral flow field within hollow cathodes. Because cathode ignition occurs typically between the keeper and orifice plates, we focused our study on the keeper-orifice region. Using the DSMC method, we showed that the neutral cathode flow results in most cases in an underexpanded, supersonic, rarefied jet with Knudsen numbers above 0.1, along with a toroidal subsonic recirculation zone. The recirculation zone is bounded by the underexpanded jet viscous layer, the cathode orifice plate, and the keeper plate. We have shown that simple conservation laws can explain the experimentally and numerically observed variation of the ratio between the keeper-orifice region pressure and the total, upstream pressure. Good agreement is obtained with our control-volume based approach and numerical results. The statistical analysis we conducted on the numerical data demonstrated that (i) the pressure within the keeper-orifice region is most likely 4% of the total, upstream pressure, although cases for which the flow remains subsonic within the keeper-orifice region result in much higher pressure recovery values (up to 80%), and (ii) $1/\beta$, the ratio between the cathode pressure-diameter product, $P_{T}d_{c}$, and the keeper-orifice region pressure-distance ratio, $P_{ko}D_{ko}$, is most likely $\approx 40$ and is related to the ratio of ignition to nominal mass flow rates.
Our study is limited to that of the neutral, noble gas, flow in cathodes and does not consider any plasma effects (e.g., neutral heating due to charge-exchange collisions) that may alter the flow structure. However, the neutral flow field obtained from DSMC may be used as a first guess in a coupled plasma / neutral numerical model (e.g., PIC/DSMC). The study is further limited by the data upon which it is based: we considered a wide, albeit non-exhaustive, range of rescaled parameters describing orificed cathodes with an enclosed keeper, and our conclusions may not apply to cathode configurations that are not covered by our dataset.
The study demonstrates that the flow within the orifice, prior to cathode ignition, is more accurately represented by a Fanno flow approach than an isentropic flow or Poiseuille flow one. The Poiseuille flow model should not be used to estimate the pressure (and/or density and temperature) within hollow cathodes because most assumptions upon which the it is based are invalid in the regime in which cathodes operate. Finally, as shown through our statistical analysis and explained by simple conservation laws, a fifty-fold increase in mass flow rate is likely to result in a minimum voltage in the Paschen-like, DC discharge. Although we only considered xenon, these conclusions are likely applicable to other noble, monatomic gases such as argon.
Acknowledgments
This work was conducted at the Princeton Collaborative Research Facility supported by the U.S. Department of Energy through contract DE-AC02-09CH11466. The computational resources of the United States Naval Academy are gratefully acknowledged.
Author declaration
Data availability
The data that support the findings of this study are available from the corresponding author upon reasonable request.
Conflict of interest
The authors have no conflicts to disclose.
Appendix
V.1 Cathode dimensions
The dimensions of the considered cathodes are shown in Table 4. The Southampton and Harbin Institute of Technology cathodes, MasMI LUC, JPL 50 A HHC are heaterless hollow cathodes. The corresponding range of normalized parameters is:
  • $L_{o}/d_{o}$: 1.1–1.5
  • $d_{o}/d_{c}$: 0.1–0.8 (excluding tube cathodes)
  • $D_{ko}/d_{o}$: 0.31–28
  • $d_{k}/d_{o}$: 1–7
Name Insert Orifice Keeper Refs.
diameter diameter length orifice diam. distance thickness
$d_{c}$ $d_{o}$ $L_{o}$ $d_{k}$ $D_{ko}$ $t_{k}$
Siegfried 3.9 0.5–1 3.6 2.5 Siegfried,1982
NEXIS 12.0 2.75 1.75 4.8 1, 2.4 1.3 Polk et al.,1999; Mikellides, Katz, and Goebel,2004
NSTAR 3.8 1.02 1.52 4.65 1 1.5 Goebel, Watkins, and Jameson,2007a; Polk et al.,1999
JPL 1.5 cm 7.0 3–5 6.4 Goebel, Watkins, and Jameson,2007b
Southampton HHC 2 0.75–2 2–7.5 1–21 2 Daykin-Iliopoulos et al.,2015
MasMI LUC 1.82 0.7 0.8 1.5 0.4 3 Becatti, Conversano, and Goebel,2021
JPL 50 A HHC 6.3 4.8 5.7 1.5 Payman and Goebel,2022
Harbin Institute 2 0.8 1.5 2.5 1.5 2 Ning et al.,2019
Table 4: Dimensions of orificed hollow cathodes with enclosed keeper from the literature. Values are in mm.
V.2 Keeper-orifice region conservation laws
We derive here Equation 15 following a framework similar to that of Emmert et al.Emmert, Lafon, and Bailly (2009) The control volume considered that is comprised of surfaces I, II, and III is shown in Figure 27. We assume that:
  1. 1.
    The flow is in the continuum regime
  2. 2.
    The flow is adiabatic
  3. 3.
    The static pressure on the walls adjacent to surfaces I, II, and III is constant and equal to the keeper-orifice region pressure, $P_{ko}$.
The subscripts $o$ and $k$ will be used to denote the quantities defined at the cathode orifice and keeper entrance surfaces, respectively.
V.2.1 Continuity
From the conservation of mass applied to the CV,
$\rho_{o}u_{o}\pi r_{o}^{2}=\rho_{k}u_{k}\pi r_{k}^{2}.$
Using the perfect gas law, $P=\rho R_{g}T$, and the definition of the Mach number, $M=u/\sqrt{\gamma R_{g}T}$, we obtain
$\dfrac{P_{o}}{P_{k}}=\dfrac{M_{k}}{M_{o}}\sqrt{\dfrac{T_{o}}{T_{k}}}\bar{d}^{2},$ (20)
where $\bar{d}=d_{k}/d_{o}$ and all thermodynamic quantities are static (as opposed to stagnation) quantities.
V.2.2 Energy
Because the flow is adiabatic, conservation of energy applied to the CV yields
$h_{T,o}=h_{T,k},$
where $h_{T}$ is the total enthalpy. Using the definition of the total enthalpy,
$h_{T}=\dfrac{\gamma}{\gamma-1}R_{g}\left[1+\dfrac{\gamma-1}{2}M^{2}\right],$
we obtain
$\dfrac{T_{o}}{T_{k}}=\dfrac{1+\dfrac{\gamma-1}{2}M_{k}^{2}}{1+\dfrac{\gamma-1}{2}M_{o}^{2}}.$ (21)
V.2.3 Momentum along $z$-axis
The conservation of momentum along the $z$-axis is
$\begin{split}-\left[\left(\rho_{o}u_{o}^{2}+P_{o}\right)\pi r_{o}^{2}+P_{ko}\pi\left(r_{e}^{2}-r_{o}^{2}\right)\right]\\ +\left[\left(\rho_{k}u_{k}^{2}+P_{k}\right)\pi r_{k}^{2}+P_{ko}\pi\left(r_{e}^{2}-r_{k}^{2}\right)\right]&\\ +\Pi_{rz}2\pi D_{ko}r_{e}&=0.\end{split}$ (22)
Using the definition of the Mach number, we obtain, after multiplying by $1/\left(P_{T}r_{o}^{2}\right)$ and simplifications:
$\begin{split}\dfrac{P_{ko}}{P_{T}}\left(\bar{d}^{2}-1\right)=\dfrac{P_{k}}{P_{o}}\cdot\dfrac{P_{o}}{P_{T}}\bar{d}^{2}\left(1+\gamma M_{k}^{2}\right)&\\ -\dfrac{P_{o}}{P_{T}}\bar{d}^{2}\left(1+\gamma M_{o}^{2}\right)&+\dfrac{\left<\tau\right>2D_{ko}r_{e}}{P_{T}ro^{2}}.\end{split}$ (23)
Inserting Equations 20 and 21 into Equation 23, and, because $r_{e}\approx 2r_{k}$, we have
$\dfrac{P_{ko}}{P_{T}}\left(\bar{d}^{2}-1\right)=\dfrac{P_{o}}{P_{T}}\mathcal{F}\left(M_{o},M_{k}\right)+8\dfrac{\left<\tau\right>}{P_{T}}\bar{d}\dfrac{D_{ko}}{d_{o}},$ (24)
where $\mathcal{F}$ is given in Equation 16.
Figure 27: Control volume considered.

### v2

@@
-Rarefied neutral flow in orificed hollow cathodes
+Rarefied xenon flow in orificed hollow cathodes
 Abstract
-A parametric study is conducted to quantify the effect of the keeper electrode geometry on the neutral flow quantities (pressure, density, temperature) within orificed hollow cathodes, prior to cathode ignition.
-The keeper impinges directly on the flow out of the cathode orifice and its geometry strongly influences the product between the pressure in the orifice-keeper region and the distance between cathode and keeper, $P_{ko}\cdot D_{ko}$, a key parameter for successful cathode ignition.
-A representative cathode equipped with a keeper is simulated using the Direct Simulation Monte Carlo method.
-The numerical model is first validated with computational results from the literature, and a parametric study is then conducted.
-Parameters include the cathode pressure-diameter in the range of 1–5 Torr-cm and the following geometric ratios: cathode orifice-to-cathode inner radii in the range of 0.1–0.7, keeper orifice-to-cathode orifice radii in the range of 1–5, and keeper distance-to-cathode-orifice diameter, in the range of 0.5–10.
-It is found that, if both keeper and cathode have identical orifice radii, the flow remains subsonic in the cathode orifice-to-keeper region.
-In most cases, however, the flow becomes underexpanded and supersonic, and the static pressure within the cathode orifice-keeper region is, on average, 4% that of the upstream pressure value.
-The orifice-keeper region pressure increases with either (i) a decrease in the keeper orifice diameter or (ii) an increase in the distance between the cathode orifice plate and the keeper plate, in agreement with literature experimental data.
-Both trends are explained through control-volume-based conservation laws.
-The ratio of ignition-to-nominal mass flow rates is found to be in the range of 9–120, with a most probable value of 40, in agreement with experimental literature data.
-This suggests that heater-less cathode ignition at a minimum DC voltage may be achieved by increasing the input mass flow rate by a factor of 40.
-The views expressed in this article are those of the authors and do not reflect the official policy or position of the Department of Defense, the U.S. Government, or any agency thereof.
-The following article has been submitted to the Journal of Applied Physics. After it is published, it will be found at https://pubs.aip.org/aip/jap.
+A parametric study is conducted to quantify the effect of the keeper electrode geometry on the xenon neutral flow quantities within orificed hollow cathodes, prior to ignition.
+The keeper impinges directly on the flow out of the cathode orifice and its geometry influences the product between the pressure in the orifice-keeper region and the cathode-to-keeper distance.
+A representative cathode is simulated using the Direct Simulation Monte Carlo method.
+The numerical model is first validated with computational results from the literature.
+A parametric study is then conducted.
+Parameters include the cathode pressure-diameter in the range of 1–5 Torr-cm and the following geometric ratios (and ranges): cathode orifice-to-inner radii (0.1–0.7), keeper orifice-to-cathode orifice radii (1–5), and keeper distance-to-cathode-orifice diameter (0.5–10).
+It is found that, if both keeper and cathode have identical orifice radii, the flow remains subsonic in the orifice-to-keeper region.
+In most cases, however, the flow becomes underexpanded and supersonic, and the static pressure within the orifice-to-keeper region is, on average, 4% that of the upstream pressure value.
+The orifice-keeper region pressure increases with either a decrease in the keeper orifice diameter or an increase in the distance between cathode and keeper, in agreement with literature data.
+Both trends are explained through conservation laws.
+A statistical study of numerical results reveals that the ratio of ignition-to-nominal mass flow rates has a most probable value of 50, which suggests that heaterless cathode ignition at a minimum DC voltage may be achieved by increasing the input mass flow rate by a factor of 50.
+The following article has been submitted to the Journal of Applied Physics.
 Copyright (2024) Pierre-Yves C.
@@
 Gas is inserted in the cathode tube, and a glow discharge produced by a high-voltage ($>100$ V) DC source effectively heats the cathode orifice and emitter.
-This approach is favored for low-to-medium amperage cathodes and circumvents the use of a heater: the heater and its associated power supply are oftentimes considered to be a single point-of-failure for space electric propulsion systems and require a dedicated power supply.
+This approach is favored for low-to-medium amperage cathodes and circumvents the use of a heater: the heater and its associated power supply are oftentimes considered to be a single point-of-failure for space electric propulsion systems.
 The accurate placement and sizing of the keeper is critical for a reliable DC discharge: an optimal value of the pressure-distance product, $P_{ko}D_{ko}$, results in a minimum in the applied voltage required for the glow discharge, where $P_{ko}$ is the pressure of the gas in the orifice-keeper region, and $D_{ko}$ is the (axial) distance between the keeper and the cathode orifice plate.
@@
 Using a DSMC approach, we seek to identify which geometrical parameters most significantly impact the pressure-distance product in the orifice-keeper region of cathodes, provide broadly applicable design rules, and explain, physically, the observed trends.
+We also compare DSMC results to that of a fluid solver in select cases to understand the limitations of continuum approaches applied to rarefied gas flows.
 We first describe our numerical approach and validate our model based on reported literature data in Section II.
@@
 The particles cannot transfer momentum nor heat — the flow is effectively inviscid.
-We choose to apply diffuse boundary conditions throughout, with an accommodation coefficient of 0.8, and do not consider the influence of gas-surface interactions on our solution.
-Other studiesSharipov (2004) have shown, for example, that the flow rate through an orifice is not sensitive to the value of the accommodation coefficient for vessels separated by a thin orifice.
+We choose to apply diffuse boundary conditions throughout with a choice of a constant accommodation coefficient for all solid surfaces.
+While the value of the accommodation coefficient may change appreciably depending on the surface material and roughness,Agrawal and Prabhu (2008); Selden et al.
+(2009) a sensitivity study of the solution on the accommodation coefficient is beyond the scope of this article: we do not consider the influence of gas-surface interactions on our solution and set the accommodation coefficient to a value of 0.8.
+While this value is 13% lower than the recommended value from Agrawal and Prabhu,Agrawal and Prabhu (2008) it is within the range of other reported data for xenon interaction with solid wallsSelden et al.
+(2009) and, therefore, represents a conservative value of the accommodation coefficient.
+Other studiesSharipov (2004) have shown that the flow rate through an orifice is not sensitive to the value of the accommodation coefficient for vessels separated by a thin orifice.
 Outflow boundary condition
@@
 Inflow boundary condition
-The upstream surface features a subsonic inlet with a set pressure and temperature.
+The upstream surface features a subsonic inlet with a set pressure and temperature , as opposed to a set mass flow rate.
 A varying number of particles is injected to enforce the desired pressure and temperature.
 The total number of injected particles accounts for domain particles that exit the simulation domain through the upstream surface.
+The mass flow rate for the numerical experiments, $\dot{m}_{\textrm{exp}}$, may be retrieved by computing the flux of particles that cross a plane intersecting the cathode channel.
+The mass flow rate typically depends on the orifice geometry and may differ from the theoretical values that can be computed using a typical isentropic flow approach, $\dot{m}_{th}$.
+We define the discharge coefficient, $C_{d}$, as the ratio of the experimental-to-theoretical flow rates:
+$C_{d}=\dfrac{\dot{m}_{\textrm{exp}}}{\dot{m}_{th}}.$ (1)
 Figure 1: Cathode computational representation.
 II.3.2 Collision model
-Particle-particle collisions are represented by the variable soft sphere (VSS) model. The corresponding VSS parameters are shown in Table 1 for xenon.
-Table 1: VSS parameters for xenon (Ref. Bird, 1994, pp. 408,410)
-Species $d_{\textrm{ref}}$(m) $\omega$ $T_{\textrm{ref}}$ $\alpha$
+Particle-particle collisions are represented by the variable soft sphere (VSS) model, in which the relationship between the collision cross-section and the collision energy is represented by a power law.
+For this model, the collision deflection angle is $\chi=2\arccos((b/d)^{\left(1/\alpha\right)})$, where $b$ is the impact parameter.
+Consequently, the gas viscosity depends on the temperature through a power law: $\mu\propto T^{\omega}$.
+The model depends on four parameters:
+  • $T_{\textrm{ref}}$: a reference temperature,
+  • $d_{\textrm{ref}}$: the particle diameter for the collision cross-section at the reference temperature, $T_{\textrm{ref}}$,
+  • $\omega$: the temperature dependence of the viscosity, and
+  • $\alpha$: the scattering parameter.
+The VSS parameters are shown in Table 1 for xenon.
+Table 1: VSS parameters for xenon (Ref.
+Bird, 1994, pp.
+408,410).
+$T_{\textrm{ref}}$, $d_{\textrm{ref}}$, $\omega$ and $\alpha$ are the reference temperature, the particle diameter for the collision cross section at the reference temperature ($T_{\textrm{ref}}$), the temperature dependence of the viscosity, and the scattering parameter, respectively.
+Species $d_{\textrm{ref}}$ (m) $\omega$ $T_{\textrm{ref}}$ (K) $\alpha$
 Xenon $5.65\times 10^{-10}$ 0.85 273.15 1.44
@@
 […]
+Figure 3: Typical evolution of the total number of particles.
+Figure 4: Illustrative adapted grid for the case where $d_{k}/d_{o}=1$ and $D_{ko}/d_{o}=0.5$.
+The cathode surface is indicated in gray.
+The internal grid is 32$\times$ more refined than the plume grid.
 Particle weight
@@
 Particles are drawn from a Maxwellian distribution with zero streaming velocity and the upstream temperature.
+After initialization, additional particles are injected continuously at the inflow boundary described in Section II.C.1.c at a set temperature and pressure.
+Injected particles travel throughout the domain and interact with the different boundary conditions.
+They are removed from the domain if they reach the outflow (or the inflow) boundary.
 The simulation is considered to have converged once the total number of particles remains approximately constant.
@@
 Because the initial density in the cathode is much lower than the steady-state one (approximately two-to-three orders of magnitude) the total number of particles increases until it reaches steady-state.
-Figure 3: Typical evolution of the total number of particles.
-Relevant physical quantities are computed once steady-state is reached. Results are averaged over 10,000 time steps ($\sim$ 0.1 ms) to decrease the statistical noise inherent to particle simulations.
+Relevant physical quantities (e.g., the gas temperature or the mass flow rate) are computed once steady-state is reached.
+Results are averaged over 10,000 time steps ($\sim$ 0.1 ms) to decrease the statistical noise inherent to particle simulations.
 II.4 Multi-level grid and load rebalance
@@
 Because the particle weight, $f_{n}$, is set globally for the simulation, a refined grid in this region would result in a low number of DSMC particles per computational cell.
-Figure 4: Illustrative adapted grid for the case where $d_{k}/d_{o}=1$ and $D_{ko}/d_{o}=0.5$.
-The cathode surface is indicated in gray.
-The internal grid is 32$\times$ more refined than the plume grid.
 Because the number of DSMC particles increases significantly within the first 10,000 steps of the simulation (see, e.g., Figure 3), we rebalance the particle load across processors every 100 steps during that time period.
 Periodic rebalancing is applied throughout the simulation thereafter.
+Figure 5: Cathode computational representation for the numerical validation test case.
 II.5 Numerical validation
@@
 (2019) which features a computational domain without a keeper electrode.
+The computational domain for this test case is shown in Figure 5.
 A fully coupled PIC/DSMC approach wherein both charged and neutral particles are simulated is beyond the scope of this article: we cannot capture plasma-neutral effects such as the increase in neutral particles temperature near the orifice due to the frequent charge-exchange collisions with the plasma ions.
 The parameters describing the computational domain of the validation case are shown in Table 2.
-Using the perfect gas law and the provided upstream pressure and density (450 Pa and 1.3$\times$1022 m-3, respectively), the temperature of the injected gas is estimated to be 2,500 K.
+Using the input parameters provided by Cao et al.Cao et al.
+(2019) for the upstream pressure and density (450 Pa and 1.3$\times$1022 m-3, respectively) and the perfect gas law, the temperature of the injected gas is found to be 2,500 K.
+Because we are considering a heavy, noble gas that is monatomic at high temperatures and low densities, the thermal de Broglie wavelength ($\mathcal{O}\left(10^{-12}\,\textrm{m}\right)$) is much smaller than the size of the domain ($\mathcal{O}\left(10^{-3}\,\textrm{m}\right)$): Boltzmann statistics still apply,McQuarrie (2000) and the perfect gas law is applicable.
 Unlike Cao et al., we do not rescale the heavy particle mass or domain size.
@@
 […]
-$\dfrac{P_{\textrm{T,exit}}}{P_{\textrm{T,in}}}=\dfrac{1}{M_{in}}\left[\dfrac{2+\left(\gamma-1\right)M_{in}^{2}}{\gamma+1}\right]^{\dfrac{\gamma+1}{2\left(\gamma-1\right)}},$ (9)
+$\dfrac{P_{\textrm{T,in}}}{P_{\textrm{T,exit}}}=\dfrac{1}{M_{in}}\left[\dfrac{2+\left(\gamma-1\right)M_{in}^{2}}{\gamma+1}\right]^{\dfrac{\gamma+1}{2\left(\gamma-1\right)}},$ (10)
 where the subscripts ${in}$ and ${exit}$ denote the orifice entrance and exit quantities, respectively.
@@
 (2023) We use the model from Goebel and KatzGoebel and Katz (2008) with the xenon viscosity from Stiel and ThodosStiel and Thodos (1961).
-The mass flow rate may be estimated by assuming a continuum flow and a choked orifice,
-$\dot{m}=C_{d}\dfrac{P_{T}\pi r_{o}^{2}}{\sqrt{T_{T}}}\sqrt{\dfrac{\gamma}{R_{g}}}\left(\dfrac{2}{\gamma+1}\right)^{\left(\gamma+1\right)/\left(\gamma-1\right)}$ (10)
-where $C_{d}$ is the discharge coefficient and $R_{g}$ the gas constant.
+The mass flow rate is computed from the simulation, as described in Section II.C.1.c, and is found to be 1.5 mg/s (15.3 sccm).
 Figure 6 compares the pressure and the neutral density as obtained with our theoretical and numerical approaches to that of Cao et al.
@@
 They are equal to 1,823 K, 447 m/s, and $5.0\times 10^{21}$ m-3, respectively.
+We estimate the average orifice gas temperature for the Poiseuille flow approach by assuming that the orifice is choked (i.e., the Mach number in the orifice is equal to one).
+The orifice static temperature used for the Poiseuille flow model is found to be equal to 1,873 K.
 Reasonable agreement is obtained for all quantities for all numerical approaches.
@@
 The Fanno flow model of the orifice yields the most accurate theoretical value.
-The Poiseuille flow approach overestimates the pressure and is, in general, inaccurate.
+The Poiseuille flow approach overestimates the pressure by 32% and 134% at the orifice entrance and exit, respectively.
 As discussed elsewhere,Taunay, Wordingham, and Choueiri (2022a) this approach should not be used to estimate the pressure within cathodes as it is inapplicable to the flow regime in which cathodes operate: most of the assumptions required by this flow model are invalid in cathodes.
@@
 The isentropic and Fanno flow models are assumed to have the same static quantities at the orifice entrance.
-The error bars on the Poiseuille flow model indicate values of $C_{d}$ in the range of 0.1–1, with the symbol set at the pressure and density values for which $C_{d}=0.5$.
 Figure 7(a) shows the neutral density contour for this test case.
@@
 […]
-Figure 6: Contours of (a) neutral density and (b) Knudsen number (normalized by the orifice diameter) for Cao et al.’s test case.
+Figure 7: Contours of (a) neutral density and (b) Knudsen number (normalized by the orifice diameter) for Cao et al.’s test case as obtained with our DSMC approach.
 A keeper electrode located at $\sim 1$ mm downstream of the orifice would impinge on the flow and affect local, static quantities.
@@
 Neutral species mass amu $M_{a}$ 131.293
+The mass flow rate may be estimated by assuming a continuum flow and a choked orifice,
+$\dot{m}=C_{d}\dfrac{P_{T}\pi r_{o}^{2}}{\sqrt{T_{T}}}\sqrt{\dfrac{\gamma}{R_{g}}}\left(\dfrac{2}{\gamma+1}\right)^{\left(\gamma+1\right)/\left(\gamma-1\right)}$ (11)
+where $C_{d}$ is the discharge coefficient and $R_{g}$ the gas constant.
 Using Equation 11 and assuming $C_{d}=1$, the range of mass flow rates considered is approximately 0.28–13.5 mg/s (2.8–135 sccm of xenon).
@@
 After reaching the keeper, the flow undergoes a second expansion, into vacuum.
-Our numerical results reveal the existence of two main flow regimes: the flow may be either supersonic or subsonic in the orifice-keeper region.
+Our numerical results reveal the existence of two main flow regimes: the flow may be either supersonic or subsonic in the orifice-keeper region, depending on the value of keeper-to-total pressure ratio, $P_{ko}/P_{T}$.
+The “keeper pressure”, $P_{ko}$, is defined as the average (static) pressure in the region bounded by the orifice plate and keeper entrance.
+Only the quiescent part of the flow (i.e., $v\approx 0$ m/s) is considered for the computation of $P_{ko}$.
+$P_{T}$ is the total upstream pressure.
+The pressure ratio we define is similar to the “ambient-to-total” (or “reservoir-to-total”) pressure ratio that is typically used to characterize nozzle flows.
 IV.1.1 Supersonic flow
 In most cases, the pressure ratio $P_{ko}/P_{T}$ is such that the flow that leaves the cathode orifice is choked and supersonic, and the overall flow structure is similar to that reported by Titarev and ShakovTitarev and Shakhov (2014) for a composite pipe system composed of two pipes of increasing radii that exhaust into vacuum.
-Figures 11 through 14 illustrate the typical flow structure in the case where the orifice is choked.
+Figures 12 through 15 illustrate the typical flow structure in the case where the orifice is choked.
 The flow expansion from the cathode orifice to the keeper-orifice region results in an underexpanded, rarefied, supersonic jet (Figures 121315).
@@
 The Knudsen number in the keeper region is $\approx 0.1$ (Figure 14), indicating that the flow is in the transitional regime.
-Refer to captionOrifice
+[…]
 Figure 12: Velocity magnitude with streamlines for the case $P_{T}d_{c}=1$ Torr-cm, $d_{o}/d_{c}=0.3$, $D_{ko}/d_{o}=3.0$, and $d_{k}/d_{o}=2.0$.
-Refer to captionOrifice
+[…]
 Figure 13: Mach number for the case $P_{T}d_{c}=1$ Torr-cm, $d_{o}/d_{c}=0.3$, $D_{ko}/d_{o}=3.0$, and $d_{k}/d_{o}=2.0$.
@@
 Figure 14: Knudsen number normalized by orifice diameter for the case $P_{T}d_{c}=1$ Torr-cm, $d_{o}/d_{c}=0.3$, $D_{ko}/d_{o}=3.0$, and $d_{k}/d_{o}=2.0$.
-Refer to captionOrifice
+[…]
 Figure 15: Static pressure for the case $P_{T}d_{c}=1$ Torr-cm, $d_{o}/d_{c}=0.3$, $D_{ko}/d_{o}=3.0$, and $d_{k}/d_{o}=2.0$.
@@
 The exact structure of the underexpanded jet depends on the backing-to-upstream pressure ratio, and, therefore, the cathode geometry.
-The centerline Mach number, static pressure, and number density are shown in Figures 16 through 19 for $d_{o}/d_{c}=$ 0.1–0.7, respectively.
+The centerline Mach number, static pressure, and number density are shown in Figures 18 through 21 for $d_{o}/d_{c}=$ 0.1–0.7, respectively.
 We observe the following qualitative trends:
@@
 This approach is beyond the scope of this work.
+A statistical distribution of the discharge coefficient as obtained with the method outlined in Section II.C.1.c is shown in Figure 17 for the supersonic flow cases.
+The mass flow rate as obtained for the isentropic mass flow rate, $\dot{m}_{th}$, that is required to compute $C_{d}$ is given by
+$\dot{m}_{th}=\dfrac{P_{T}\pi r_{o}^{2}}{\sqrt{T_{T}}}\sqrt{\dfrac{\gamma}{R_{g}}}\left(\dfrac{2}{\gamma+1}\right)^{\left(\gamma+1\right)/\left(\gamma-1\right)}.$ (14)
+The average value and 95% confidence bounds for $C_{d}$ are 0.82, and 0.55 and 0.92, respectively.
+The study of the variation of $C_{d}$ with the geometric parameters (i.e., $d_{o}/d_{c}$, $d_{k}/d_{o}$, and $D_{ko}/d_{o}$) is beyond the scope of this article.
+For a cathode with a similar geometry as the one studied and for a set upstream pressure, a range of mass flow rate may be estimated with Equation 11 and $C_{d}$ in the 95% confidence interval.
+This approach may also be used to estimate the upstream (total) pressure of the gas, $P_{T}$, for a given mass flow rate.
+Figure 17: Combined violin and box plot of the range of discharge coefficient, $C_{d}$. Each individual tick mark indicates the location of a data point.
 IV.1.2 Subsonic flow
@@
 Because the flow is subsonic in this particular case, the keeper-to-total pressure ratio, $P_{ko}/P_{T}$, is higher than that in the supersonic case.
-Figures 20 and 21 show a contour of the Mach number and static pressure for the subsonic case, respectively.
+Figures 22 and 23 show a contour of the Mach number and static pressure for the subsonic case, respectively.
 For high-enough mass flow rates, however, the flow may be supersonic.
 The exact point at which the flow transitions from subsonic to supersonic for the case where $d_{o}=d_{k}$ is beyond the scope of this article.
+Some heaterless cathodes may have a configuration for which $d_{k}<d_{o}$ (i.e., $d_{k}/d_{o}<1$).
+This configuration is beyond the scope of this article: a continuum or near-continuum solver is likely more appropriate for this case as the flow will remain subsonic and in the continuum regime throughout the orifice-keeper region.
+This case is illustrated in Figure 24, which shows the total pressure and Mach number for the case where $d_{k}/d_{o}=0.5$, with $d_{o}/d_{c}=0.3$ and $D_{ko}/d_{o}=3$.
+The fluid solver (ANSYS-CFX) was used for this particular case.
+The total pressure remains approximately constant ($\approx 1$ Torr) in the cathode channel and the orifice-keeper region.
+Much like the case $d_{o}=d_{k}$, the supersonic expansion occurs in the keeper orifice.
 Figure 18: Centerline values for $d_{o}/d_{c}=0.1$ and $P_{T}d_{c}=1$ Torr-cm.
@@
 The horizontal axis is normalized by the orifice-keeper distance, $D_{ko}$.
+A scaled schematic of each configuration investigated is shown under each $d_{k}/d_{o}$ case.
+The keeper entrance for $D_{ko}/d_{o}=1$ overlaps the keeper exit for $D_{ko}/d_{o}=0.5$
 Figure 19: Centerline values for $d_{o}/d_{c}=0.3$ and $P_{T}d_{c}=1$ Torr-cm.
@@
 The horizontal axis is normalized by the orifice-keeper distance, $D_{ko}$.
+A scaled schematic of each configuration investigated is shown under each $d_{k}/d_{o}$ case.
+The keeper entrance for $D_{ko}/d_{o}=1$ overlaps the keeper exit for $D_{ko}/d_{o}=0.5$
 Figure 20: Centerline values for $d_{o}/d_{c}=0.5$ and $P_{T}d_{c}=1$ Torr-cm.
@@
 The horizontal axis is normalized by the orifice-keeper distance, $D_{ko}$.
+A scaled schematic of each configuration investigated is shown under each $d_{k}/d_{o}$ case.
+The keeper entrance for $D_{ko}/d_{o}=1$ overlaps the keeper exit for $D_{ko}/d_{o}=0.5$
 Figure 21: Centerline values for $d_{o}/d_{c}=0.7$ and $P_{T}d_{c}=1$ Torr-cm.
@@
 The horizontal axis is normalized by the orifice-keeper distance, $D_{ko}$.
+A scaled schematic of each configuration investigated is shown under each $d_{k}/d_{o}$ case.
+The keeper entrance for $D_{ko}/d_{o}=1$ overlaps the keeper exit for $D_{ko}/d_{o}=0.5$
 Figure 22: Mach number for the (subsonic) case $P_{T}d_{c}=1$ Torr-cm, $d_{o}/d_{c}=0.3$, $D_{ko}/d_{o}=3.0$, and $d_{k}/d_{o}=1.0$.
 Figure 23: Static pressure for the (subsonic) case $P_{T}d_{c}=1$ Torr-cm, $d_{o}/d_{c}=0.3$, $D_{ko}/d_{o}=3.0$, and $d_{k}/d_{o}=1.0$.
+Figure 24: (a) Total pressure and (b) Mach number for the case $d_{o}/d_{c}=0.3$, $D_{ko}/d_{o}=3$, and $d_{k}/d_{o}=0.5$.
 IV.2 Keeper pressure-distance product
@@
 Good agreement is obtained with our control-volume based approach and numerical results.
-The statistical analysis we conducted on the numerical data demonstrated that (i) the pressure within the keeper-orifice region is most likely 4% of the total, upstream pressure, although cases for which the flow remains subsonic within the keeper-orifice region result in much higher pressure recovery values (up to 80%), and (ii) $1/\beta$, the ratio between the cathode pressure-diameter product, $P_{T}d_{c}$, and the keeper-orifice region pressure-distance ratio, $P_{ko}D_{ko}$, is most likely $\approx 40$ and is related to the ratio of ignition to nominal mass flow rates.
-Our study is limited to that of the neutral, noble gas, flow in cathodes and does not consider any plasma effects (e.g., neutral heating due to charge-exchange collisions) that may alter the flow structure.
+The statistical analysis we conducted on the numerical data demonstrated that (i) the pressure within the keeper-orifice region is most likely 4% of the total, upstream pressure, although cases for which the flow remains subsonic within the keeper-orifice region result in much higher pressure recovery values (up to 80%), and (ii) $1/\beta$, the ratio between the cathode pressure-diameter product, $P_{T}d_{c}$, and the keeper-orifice region pressure-distance ratio, $P_{ko}D_{ko}$, is most likely $\approx 50$ (excluding the case where $d_{k}=d_{o}$) and is related to the ratio of ignition to nominal mass flow rates.
+Our study is limited to that of the neutral xenon flow in cathodes and does not consider any plasma effects (e.g., neutral heating due to charge-exchange collisions) that may alter the flow structure.
 However, the neutral flow field obtained from DSMC may be used as a first guess in a coupled plasma / neutral numerical model (e.g., PIC/DSMC).
@@
 The Poiseuille flow model should not be used to estimate the pressure (and/or density and temperature) within hollow cathodes because most assumptions upon which the it is based are invalid in the regime in which cathodes operate.
+Select numerical cases also show that fluid solvers, while offering qualitative agreement with DSMC results in the keeper-orifice region, differ quantitatively from DSMC in regions where the gas becomes transitional and rarefied.
 Finally, as shown through our statistical analysis and explained by simple conservation laws, a fifty-fold increase in mass flow rate is likely to result in a minimum voltage in the Paschen-like, DC discharge.
-Although we only considered xenon, these conclusions are likely applicable to other noble, monatomic gases such as argon.
+Although we only considered xenon, these conclusions are likely applicable to other noble, monatomic gases such as argon: while a lighter gas will have a lower mass flow rate for the same total pressure, temperature, and geometry (see Equation 11), all other scaled parameters (Mach number, Knudsen number, pressure-diameter product, ratio of specific heats, geometric parameters) remain similar.
 Acknowledgments
+The views expressed in this article are those of the authors and do not reflect the official policy or position of the U.S.
+Naval Academy, Department of the Navy, the Department of Defense, the U.S.
+Government, or any agency thereof.
 This work was conducted at the Princeton Collaborative Research Facility supported by the U.S.
@@
 Using the perfect gas law, $P=\rho R_{g}T$, and the definition of the Mach number, $M=u/\sqrt{\gamma R_{g}T}$, we obtain
-$\dfrac{P_{o}}{P_{k}}=\dfrac{M_{k}}{M_{o}}\sqrt{\dfrac{T_{o}}{T_{k}}}\bar{d}^{2},$ (20)
+$\dfrac{P_{o}}{P_{k}}=\dfrac{M_{k}}{M_{o}}\sqrt{\dfrac{T_{o}}{T_{k}}}\bar{d}^{2},$ (B1)
 where $\bar{d}=d_{k}/d_{o}$ and all thermodynamic quantities are static (as opposed to stagnation) quantities.
@@
 we obtain
-$\dfrac{T_{o}}{T_{k}}=\dfrac{1+\dfrac{\gamma-1}{2}M_{k}^{2}}{1+\dfrac{\gamma-1}{2}M_{o}^{2}}.$ (21)
+$\dfrac{T_{o}}{T_{k}}=\dfrac{1+\dfrac{\gamma-1}{2}M_{k}^{2}}{1+\dfrac{\gamma-1}{2}M_{o}^{2}}.$ (B2)
 V.2.3 Momentum along $z$-axis
 The conservation of momentum along the $z$-axis is
-$\begin{split}-\left[\left(\rho_{o}u_{o}^{2}+P_{o}\right)\pi r_{o}^{2}+P_{ko}\pi\left(r_{e}^{2}-r_{o}^{2}\right)\right]\\ +\left[\left(\rho_{k}u_{k}^{2}+P_{k}\right)\pi r_{k}^{2}+P_{ko}\pi\left(r_{e}^{2}-r_{k}^{2}\right)\right]&\\ +\Pi_{rz}2\pi D_{ko}r_{e}&=0.\end{split}$ (22)
+$\begin{split}-\left[\left(\rho_{o}u_{o}^{2}+P_{o}\right)\pi r_{o}^{2}+P_{ko}\pi\left(r_{e}^{2}-r_{o}^{2}\right)\right]\\ +\left[\left(\rho_{k}u_{k}^{2}+P_{k}\right)\pi r_{k}^{2}+P_{ko}\pi\left(r_{e}^{2}-r_{k}^{2}\right)\right]&\\ +\Pi_{rz}2\pi D_{ko}r_{e}&=0.\end{split}$ (B3)
 Using the definition of the Mach number, we obtain, after multiplying by $1/\left(P_{T}r_{o}^{2}\right)$ and simplifications:
-$\begin{split}\dfrac{P_{ko}}{P_{T}}\left(\bar{d}^{2}-1\right)=\dfrac{P_{k}}{P_{o}}\cdot\dfrac{P_{o}}{P_{T}}\bar{d}^{2}\left(1+\gamma M_{k}^{2}\right)&\\ -\dfrac{P_{o}}{P_{T}}\bar{d}^{2}\left(1+\gamma M_{o}^{2}\right)&+\dfrac{\left<\tau\right>2D_{ko}r_{e}}{P_{T}ro^{2}}.\end{split}$ (23)
-Inserting Equations 20 and 21 into Equation 23, and, because $r_{e}\approx 2r_{k}$, we have
-$\dfrac{P_{ko}}{P_{T}}\left(\bar{d}^{2}-1\right)=\dfrac{P_{o}}{P_{T}}\mathcal{F}\left(M_{o},M_{k}\right)+8\dfrac{\left<\tau\right>}{P_{T}}\bar{d}\dfrac{D_{ko}}{d_{o}},$ (24)
+$\begin{split}\dfrac{P_{ko}}{P_{T}}\left(\bar{d}^{2}-1\right)=\dfrac{P_{k}}{P_{o}}\cdot\dfrac{P_{o}}{P_{T}}\bar{d}^{2}\left(1+\gamma M_{k}^{2}\right)&\\ -\dfrac{P_{o}}{P_{T}}\bar{d}^{2}\left(1+\gamma M_{o}^{2}\right)&+\dfrac{\left<\tau\right>2D_{ko}r_{e}}{P_{T}ro^{2}}.\end{split}$ (B4)
+Inserting Equations B1 and B2 into Equation B4, and, because $r_{e}\approx 2r_{k}$, we have
+$\dfrac{P_{ko}}{P_{T}}\left(\bar{d}^{2}-1\right)=\dfrac{P_{o}}{P_{T}}\mathcal{F}\left(M_{o},M_{k}\right)+8\dfrac{\left<\tau\right>}{P_{T}}\bar{d}\dfrac{D_{ko}}{d_{o}},$ (B5)
 where $\mathcal{F}$ is given in Equation 18.